### 16 Feb 2024 v2

# Any-point Trajectory Modeling for Policy Learning

Abstract
[…]
Figure 1: Given a task instruction and the initial positions of any set of points in an image frame, our Any-point Trajectory Model (ATM) can predict the future trajectories of these points conditioned on the task. After training the model on a video dataset, the predicted trajectories serve as effective guidance for learning visuomotor policies for a set of language-conditioned manipulation tasks.
footnotetext: *First three authors contributed equally: Chuan Wen led the implementation and experiments. Xingyu Lin came up with the idea, supervised the technical development, and contributed to model debugging. John So implemented the Robot-to-robot transfer experiments and UniPi baselines.
I Introduction
Computer vision and natural language understanding have recently made significant advances [20, 7], largely due to the availability of large datasets. Similarly, in robotics, scaling up human demonstration data has been key for learning new skills [6, 32, 13], with a clear trend of performance improvement with larger datasets [27, 6]. However, human demonstrations, typically action-labeled trajectories collected via teleoperation devices [52, 49], are time-consuming and labor-intensive to collect. For instance, collecting $130K$ trajectories in [6] took 17 months, making data collection a major bottleneck in robot learning.
Videos contain knowledge about behaviors, physics, and semantics, presenting an alternative data source. However, the lack of action labels makes utilization of video data in policy learning difficult. Previous works have addressed this by using self-supervised objectives for video pre-training to learn a feature representation of the observation for policy learning [41, 31, 39]. However, a feature representation only describes the state at the current time step, largely neglecting the transition dynamics that predicts future states. To explicitly model the transition dynamics, prior works have developed video prediction models that predict future image frames from current ones [11, 51, 12]. However, learning a video prediction model for control introduces two issues. Firstly, the task of video prediction avoids any abstraction by modeling changes to every pixel, coupling the physical motion with visual appearances such as texture, and lighting. This coupling makes modeling difficult, often resulting in hallucinations and unrealistic future predictions [11]. Secondly, these models are computationally demanding in both training and inferencing. With limited computational resources, performance significantly declines. Moreover, the high inference cost compels these models to adopt open-loop execution [11, 5], which tends to result in less robust policies.
In this paper, we propose a novel and structured representation to bridge video pre-training and policy learning. We first represent each state as a set of points in a video frame. To model the temporal structure in videos, we learn an Any-point Trajectory Model (ATM) that takes the positions of the points in the current frame as input and outputs their future trajectories. We predict these trajectories in the camera coordinate frame to minimize the assumptions on calibrated cameras. These 2D point trajectories correspond to trajectories of particles in the 3D space and are a universal representation of the motions that can transfer to different domains and tasks. Contrasting with the video prediction approach of tracking changes in pixel intensity, our particle-based trajectory modeling offers a more faithful abstraction of the physical dynamics, naturally incorporating inductive biases like object permanence. We first pre-train the trajectory model on actionless video datasets. After pre-training, the predicted trajectories serve as detailed guidance for the policies, functioning like subgoals. We then train trajectory-guided policies using only a minimal amount of action-labeled demonstration data. For training the ATMs, we generate self-supervised training data by leveraging recent advancements in vision models for accurate point tracking [18]. Across over 130 language-conditioned tasks we evaluated in both simulation and the real world, ATM significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. Additionally, we demonstrate that our method facilitates effective transfer learning from human videos and videos of a robot with a different morphology. We summarize our main contributions below:
  1. 1.
    We propose an Any-point Trajectory Model, a simple and novel framework that bridges video pre-training to policy learning, leveraging the structured representation of particle trajectories.
  2. 2.
    Through extensive experiments on simulated benchmarks and in the real world, we demonstrate that our method can effectively utilize video data in pre-training and significantly outperform various video pre-training baselines in an imitation learning setting.
  3. 3.
    We demonstrate effective learning from human videos.
II Related Work
State representation for control. In learning end-to-end visuomotor policies, the policy is typically parameterized as a neural network that takes image observation as the input representation [36, 6, 46]. Due to the lack of inductive bias, these approaches require training on a large number of demonstration trajectories, which is expensive to collect. On the other hand, different structured representations are proposed to improve the data efficiency, such as key points [34, 29], mesh [23, 17], or neural 3D representation [22, 35]. However, prior structures often limit the policy to specific tasks. In contrast, we propose to utilize future trajectories of arbitrary points in the image as additional input to the policy, making minimal assumptions about the task and environment. We demonstrate its wide application to a set of over 130 language-conditioned manipulation tasks. In navigation and locomotion, it is common to construct policies that are guided by the future trajectories of the robot [1, 33]. In manipulation, some works have explored flow-based guidance [14, 38, 15]. However, prior works only track task-specific points, such as the end-effector of the robot, or the human hand. Instead, our approach works with arbitrary points, including points on the objects, thus providing richer information in the more general settings. Finally, Vecerik et al. [45] proposes to utilize any-point tracking for few-shot policy learning. This approach does not learn trajectory models from data, but instead mainly uses the tracker to perform visual servoing during test time. This design choice requires more instrumentation, such as separating the task into multiple stages, predicting the goal locations of the points for each stage, and running the tracker during inference time. In contrast, we present a much simpler framework, enabling application in more diverse settings.
Video pre-training for control. Videos contain rich information about behaviors and dynamics, which can help policy learning. However, video pre-training remains challenging due to the lack of action labels. One line of works first learns an inverse dynamics model that predicts the action from two adjacent frames and then labels the videos with pseudo actions [37, 3, 44]. However, the inverse dynamics model is often trained on a limited action-labeled dataset and does not generalize well, especially for continuous actions. Prior works have also explored pre-train a feature representation using various self-supervised objectives [40, 31, 26], but the representation alone does not retain the temporal information in videos. More recently, learning video prediction as pre-training has shown promising results [11, 21, 4, 51, 12]. During policy learning, a video prediction model is used to generate future subgoals and then a goal-conditioned policy can be learned to reach the sub-goal. However, video prediction models often result in hallucinations and unrealistic physical motions. Furthermore, video models require extensive computation, which is an issue, especially during inference time. In contrast, our method models the trajectories of the points, naturally incorporating inductive bias like object permanence while requiring much less computation. This enables our trajectory models to be run closed-loop during policy execution. Furthermore, the trajectories provide dense guidance to the policy as a motion prior.
Learning from Human Videos. Of particular interest, numerous prior works have proposed learning from the rich source of human videos [50, 42, 30, 2, 43]. However, these works often explicitly extract the hand pose or contact regions from the human videos, thereby losing information about the dynamics of the remaining objects. In contrast, our method models the trajectories of arbitrary points and can learn from both human videos and videos of a different robot.
Figure 2: Overview of our framework. (a) In the first stage, given an action-free video dataset, we first sample 2D points on one video frame and track their trajectories throughout the video using a pre-trained tracker. We then train a track transformer to predict future point trajectories given the current image observation, the language instruction, and the initial positions of the points. For the transformer input, we replace the future point positions with masked values. (2) In the second stage, we learn a track-guided policy to predict the control actions. Guidance from the predicted track enables us to learn robust policies from only a few action-labeled demonstrations.
III Preliminary
In this paper, we aim to learn robust control policies from a small set of action-labeled demonstration trajectories. Our central goal is to leverage the more scalable, unlabeled videos as a data source for pre-training.
Imitation from demos and videos. To begin with, we denote the action-free video dataset as $\mathcal{T}_{o}=\{(\tau_{o}^{(i)},\ell^{(i)})\}_{i=1}^{N_{o}},$ where $\ell^{(i)}$ is the language instruction for the $i^{th}$ episode and $\tau_{o}^{(i)}=\{o_{t}^{(i)}\}_{t=1}^{T}$ denotes the observation-only trajectory consisting of camera images. Similarly, we denote the demonstration dataset as $\mathcal{T}_{a}=\{(\tau_{a}^{(i)},\ell^{(i)})\}_{i=1}^{N_{a}},$ where $\tau_{a}^{(i)}=\{o_{t}^{(i)},a_{t}^{(i)}\}_{t=1}^{T}$ is the action-labeled trajectory. During imitation learning, our goal is to learn a policy $\pi_{\theta}$, parameterized by $\theta$ to mimic the expert behavior by the following behavioral cloning objective:
$\pi_{\theta}^{*}=\arg\min_{\theta}E_{(o_{t},a_{t},\ell)\sim\mathcal{T}_{a}}\Big{[}\mathcal{L}\Big{(}\pi_{\theta}(o_{t},\ell),a_{t}\Big{)}\Big{]},$ (1)
where $\mathcal{L}$ is the loss function, which could be Mean Squared Error (MSE) or cross-entropy loss.
Tracking Any Point (TAP). The recent advancements in video tracking [9, 10, 47] enable us to track the trajectory of each point in video frames without external supervision. In this paper, we utilize the off-the-shelf tracker proposed in [18]. Formally, given a sequence of images from a video $o_{1},...,o_{T}$, any one of the time steps $\bar{t}\in[1,T]$, and a set of points in that frame $\{p_{\bar{t},k}\}_{k=1}^{K}$, where $p_{\bar{t},k}=(x,y)$ is the point coordinate in the camera frame, the task of tracking is to predict the 2D camera-frame coordinates of the corresponding points in every frame $p_{t,k}$ where $t=1\dots T$. In this paper, we use the terms trajectory and track interchangeably to refer to a sequence of 2D coordinates of any point, denoted as $(p_{1},\dots,p_{T})$. We only model the 2D trajectories in the camera frame so that we do not have to make additional assumptions about multi-view cameras, or the availability of depth, allowing future scaling to more diverse video datasets. The tracker additionally predicts a binary visibility value $v_{t,k}$ denoting whether the point is occluded at step $t$.
IV Method
Videos contain a great deal of prior information about the world, capturing physical dynamics, human behaviors, and semantics that are invaluable for policy learning. Beyond just learning representations from videos [40, 31, 26], we aim to learn a model from videos to predict future states for guiding a control policy. In this way, we can essentially decompose the visuomotor policy learning challenge into two parts. The first part is learning what to do next by generating future states as concrete sub-goals, which is learned purely from videos. The second part is learning to predict control actions to follow the sub-goals, which require much less data to train compared to learning policies end-to-end. With sufficient video pre-training, we will be able to learn generalizable policies even from limited action-labeled trajectories. Prior works [11, 21, 5] have predominantly relied on pixel-level future frame prediction as video pre-training. While video prediction is resource-intensive during both training and inference stages, its focus on reconstructing pixel-level details, which are often extraneous to policy learning, can adversely affect the efficiency of subsequent policy learning.
We propose Any-point Trajectory Modeling (ATM). As illustrated in Figure 2, ATM learns to predict future point trajectories in a video frame as the pre-training, then uses the predicted trajectories to guide policy learning. Our proposed method will be comprehensively detailed in this section: In Sec. IV-A, we first describe how to learn a point trajectory prediction model from an action-free video dataset $\mathcal{T}_{o}$. Then in Sec. IV-B, we outline how we utilize the pre-trained track prediction model to learn a track-guided policy from a limited action-labeled trajectory datasets $\mathcal{T}_{a}$.
IV-A Trajectory Modeling from Video Datasets
Our goal is to pre-train a model from videos that forecasts the future point trajectories in a frame. More formally, given an image observation $o_{t}$ at timestep $t$, any set of 2D query points on the image frame $\mathbf{p}_{t}=\{p_{t,k}\}_{k=1}^{K}$, and a language instruction $\ell$, we learn a model $\mathbf{p}_{t:t+H}=\tau_{\theta}(o_{t},\mathbf{p}_{t},\ell)$ that predicts the coordinates of the query points in the future $H$ steps in the camera frame, where $\mathbf{p}_{t:t+H}\in\mathbb{R}^{H\times K\times 2}$. To model the tracks, we propose a track transformer and illustrate the architecture in Figure 2 (a).
Self-supervised Track Annotation.
Initially, we generate point trajectories from action-free videos for trajectory modeling pre-training. As described in Sec. III, we employ a vision tracker to pre-process videos and generate a tracking dataset. For each video, we randomly sample a time step $\bar{t}$ and then randomly sample points on this frame and generate their tracks by running the tracker. However, for a static camera, most of the points that are sampled randomly will be in the background, thus providing little information when training our track transformer. To address this, we adopt a heuristic solution to filter out these static points: we first sample a grid of $n~{}\times~{}n$ on frame $\bar{t}$ and track the grid of points across the whole of the video to obtain an initial set of tracks $\tau\in\mathbb{R}^{n^{2}\times T\times 2}$. Subsequently, we filter points that have not moved during the video by thresholding the variance of the point positions over time. In the final step, we resample points around the filtered locations and finally generate their positions using the tracker.
Multimodal Track Modeling.
We formalize the future forecasting problem as a multi-modal masked prediction problem: we aim to predict the future positions of each point, conditioned on its current position, the current image observation, and a language instruction of the task. We first encode different modalities into a shared embedding space, each represented by a few tokens. For the tracks, we mask out the future positions of all points before encoding and then separately encode each point into one token. For the language instruction, we use a pre-trained BERT [8] encoder. For the images, we split them into image patches and randomly masked out $50\%$ of the patches. We then pass all tokens through a large transformer model. Finally, we decode the track tokens into future trajectories of the corresponding points. Additionally, we reconstruct the image patches from the corresponding tokens following He et al. [16] as an auxiliary task, which we find useful for more complex tasks. Through this pre-training process, our track transformer learns the motion prior of particles within the video frames.
IV-B Track-guided Policy Learning
After training a track transformer to predict future tracks based on observations, we can then learn policies guided by these predicted trajectories.
Arbitrary Points Tracking.
During track transformer pre-training, we can filter tracks without large movements. However, using this heuristic requires knowing the future positions of each point, which can be expensive to compute during policy inference. Instead, we find it sufficient to simply use a fixed set of 32 points on a grid for the policy. This sampling method avoids the potential complexities of learning key points or finding points to track [45] and works well in practice. ATM is permutation invariant to the input set of points, and we also find ATM to be robust to the distribution of the points, allowing us to use a different point sampling scheme from training for policy learning.
Figure 3: A visual illustration of the architecture of the track-guided policy. Given the current observation and the predicted tracks from the frozen pre-trained track transformer, we train a track-guided policy from a limited demonstration dataset.
Track-guided Policy Learning.
A track-guided policy $\pi(a_{t}|o_{t},\textbf{p}_{t:t+H})$ takes input the current observation $o_{t}$ and the predicted tracks $\textbf{p}_{t:t+H}$ and predict the actions. A simplified illustration of our policy architecture is shown in Figure 3. Our transformer policy architecture follows the architecture in prior works [25, 19]. Although the predicted tracks alone already provide rich information to predict the actions, we still incorporate contextual image observations into our policy so that no information is lost, as suggested in prior works [24]. We incorporate the track tokens both before and after fusion with the image tokens (early fusion and late fusion) to ensure that the guiding information from tracks can be effectively integrated. Surprisingly, as the tracks already provide the fine-grained subgoals, we find that the policy no longer needs language instruction at this stage as task specification. Essentially, the provided tracks have reduced the difficult policy learning problems into a much easier sub-goal following problems, reducing the policy into an inverse dynamics model. Our track-guided policy is trained with MSE loss. A detailed architecture diagram and hyperparameters are available in the appendix.
V Experiments
We perform experiments to answer the following questions:
  • How does ATM compare with state-of-the-art video pre-training baselines in learning from action-free videos?
  • Can ATM be used for learning from video data that are out of the distribution of demonstration data?
  • How do various design choices of ATM affect the performance?
Our experiments are split into three sections. In Sec. V-A, we compare ATM with video pre-training baselines on over 130 language-conditioned manipulation tasks in simulation and in the real world. In Sec. V-B, we show that ATM enables effective learning from human videos. Finally, we present ablation results in Sec. V-C.
Figure 4: We compare with state-of-the-art video pre-training methods on language-conditioned manipulation tasks in the LIBERO benchmark [25]. (a) Visualization of the LIBERO tasks separated into four suites, focusing on different aspects of the manipulation policies in spatial reasoning, object reasoning, task understanding, and performing long-horizon tasks. (b) Quantitative comparisons on different suites. We additionally compare baselines with fast computation on a task suite containing 90 tasks (i.e. LIBERO-90). ATM outperforms the baselines in all tasks and excels in LIBERO-Goal and LIBERO-Long.
Figure 5: Real robot experiments on a dining table setup consisting of five tasks. The left figure shows our real-world setup and the tasks. The top right figure shows an example of the predicted particle trajectories and the policy execution, which closely follows the predicted trajectories. From the quantitative results, we can see that ATM shows significant improvements over state-of-the-art video pre-training baselines on average.
V-A Video Pre-training for Imitation Learning
Environments. We compare with baselines on over one hundred language-conditioned manipulation tasks in the LIBERO benchmark [25]. The benchmark is subdivided into five suites, LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. Each suite has 10 tasks, except LIBERO-90 which contains 90 tasks. Each task comes with expert human demonstrations. LIBERO-Spatial contains tasks with the same objects but different layouts; LIBERO-Object has tasks with the same layouts but different objects; LIBERO-Goal has tasks with the same object categories and spatial layouts, but different goals; LIBERO-Long has tasks with diverse object categories and layouts, and long-horizon task goals; LIBERO-90 has extremely diverse object categories, layouts and task goals.
Data. We compare with baselines on each suite separately. All methods are trained on $10$ action-labeled demonstration trajectories and $50$ action-free video demonstration trajectories of the robot for each task, amounting to $500$ videos for each 10-task suite. The demonstration dataset contains RGB images from a third-person camera and a wrist camera, together with gripper and joint states as observations. As each task is specified by a language instruction, we use the pre-trained BERT network to obtain a task embedding [8]. Image resolution is $128\times 128$ and the action space is 7-dimension, representing the translation, rotation, and aperture of the end-effector.
Baselines. We compare with the following baselines:
  1. 1.
    BC denotes the vanilla behavioral cloning which trains a policy exclusively using the limited action-labeled expert demonstrations, without using the video dataset. It uses a policy architecture identical to ATM except that the particle trajectories are masked to be zero and it instead takes language embedding as task specification.
  2. 2.
    R3M-finetune [31] uses a contrastive learning objective for learning representation that aligns video and language with a combination of time contrastive losses, $L_{1}$ regularization, and language consistency losses. We adopt the publicly released Ego4D pre-trained weights and fine-tune the weights on our in-domain video dataset $\mathcal{T}_{o}$, to initialize the behavioral cloning policy’s visual encoder. During policy training, we also further fine-tune the R3M backbone with the behavioral cloning loss. While this method captures priors from action-free videos through representation learning, the visual representation lacks knowledge about the transition dynamics critical to decision-making.
  3. 3.
    VPT [3] first trains an inverse dynamics model from the action-labeled trajectories $\mathcal{T}_{a}$ and then uses it to predict pseudo action labels for the video dataset. With these pseudo-labels, a policy is then trained with behavioral cloning. This method requires the inverse dynamics model to be robust to a wide distribution of input observation, which can be difficult to learn from the limited demonstrations.
  4. 4.
    UniPi [11, 21, 5] trains a text-conditioned video diffusion model to generate a temporally fine-grained video plan from an initial frame and a language instruction. During policy learning, UniPi trains an inverse dynamics model with action-labeled data. We base our implementation off of the UniPi implementation in Ko et al. [21]. While both UniPi and ATM leverage a policy conditioned on future subgoals, a trajectory representation decouples motion from other pixel-based information and makes policy learning much easier. Please see the Appendix where we perform additional comprehensive comparisons of different variations of UniPi.
Results. We present the main results in Figure 4. We see that by bridging the video data and policy learning with the structured representation of point trajectories, ATM (our method) significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. The comparison of BC with ATM shows that learning from additional videos provides useful information for policy learning. VPT performs poorly as we empirically observed that the pseudo-action labels predicted by VPT generally show large errors on the video dataset. UniPi fails on more complex as video prediction models are not physically grounded and often generate future frames that are not physically feasible, such as cases where robots disappear from the image. Please see our video for failure cases of a video prediction model.
Real World Setup. We conduct a language-conditioned manipulation experiment in the real world to further strengthen our claim. As illustrated in Figure 5, we learn policies for a 6-DOF UR5 robot arm using human expert demonstrations collected with the GELLO teleoperation system [49]. The action space is joint position control and gripper state. The observations include two RGB images from two RealSense cameras. To compensate for the partial observation, we stack the most recent two frames as the input of agents. We do not feed proprioception states into agents because it is reported that imitation policies tend to overfit them and ignore the visual inputs, leading to worse online rollout performance [28, 24]. We collect a total of $50$ action-labeled trajectories and $250$ action-free demonstration videos (by simply removing the actions from the collected action-labeled demonstrations). From the quantitative results in Figure 5, we see consistent trends that ATM outperforms BC and other video pre-training baselines by a significant margin.
Figure 6: Learning robotic skills from human videos for three tasks. We collect 100 videos of a human performing the tasks directly and 10 teleoperation demonstration trajectories. Each row from the top to the bottom shows three snapshots from the human videos, ATM trained without the human videos, and ATM trained with the human videos. By comparison, we can see that human videos are critical in learning accurate trajectory prediction and enable the policy to successfully perform the task.
Figure 7: Cross-morphology skill transfer for a pick-and-place task. Here, we collect 160 action-free videos of a Franka arm and 10 action-labeled demonstrations from a UR arm, with the final goal of learning a UR policy. We compare a vanilla BC baseline with ATM trained using types of data: using only the 10 UR videos, using only the 160 Franka videos, and using both Franka and UR videos (Franka $\Rightarrow$ UR). In the right plot, we observe that the additional cross-embodiment data led to significantly better results. Surprisingly, even if the trajectory model is only trained using Franka videos, it exhibits much better performance than the BC without the Franka videos.
TABLE I: Average success rates of human-to-robot experiments. ATM trained with human videos significantly outperforms BC and ATM trained with only 10 robot videos, demonstrating the cross-embodiment capability of ATM.
Method
Teleoperation
demos
Human
videos
fold cloth put tomato sweep toys
BC 0% 10% 30%
ATM 0% 0% 13%
ATM 63% 63% 60%
V-B Human-to-robot and Robot-to-robot Transfer
By modeling the low-level any-point trajectories, ATM enables learning from cross-embodiment videos of humans or a different robot performing the task. This facilitates the use of more scalable data sources. To verify this, we present the results of learning from human videos in Fig. 6, with the quantitative results presented in Table I, and the results of cross-robot transfer in Fig. 7. We compare the following methods: (1) Learning behavior cloning policies only on the action-labeled data (2) ATM, using a track transformer trained only on the limited action-labeled robot data, and (3) ATM, using a track transformer trained on both the action-free human and action-labeled robot data. Experiments show that training the trajectory model on additional cross-embodiment videos makes the trajectory prediction more robust and accurate, significantly improving policy learning. On the other hand, as the number of action-labeled trajectories is small, BC baselines that only use action-labeled trajectories fail. Please refer to the videos on our website for better visualization.
V-C Ablation Analysis
We conduct a series of ablation experiments in simulation to demonstrate the effect of our design choices. These include the number of action-labeled trajectories needed for policy learning, the effect of the trajectory prediction horizon, image masking, and late track fusion.
Figure 8: Success rate of our policy trained with 4%, 10% and 20% action-labeled demos. Our policy trained with only 4% demos performs comparably to BC baseline with 20% demos on LIBERO-Spatial, Object, and GOAL, and even better on LIBERO-Spatial. When trained on 20% demos, our performance approaches BC with all training data.
[…]
Figure 9: We plot the success rates of the policies learned with predicted trajectories of different lengths. Generally, longer trajectory length improves the performance, but the benefit tends to plateau at 16.
TABLE II: Ablation study on image masking of track transformer, where “w/o image masking” represents that we do not mask out image patches during track transformer training and “w/ image masking” means we randomly mask 50% patches. We can see that mask image modeling in track transformer improves the policy performance.
Image Mask Ratio Spatial Object Goal Long
w/o image masking $69.17\pm 6.38$ $65.00\pm 3.89$ $74.33\pm 3.66$ $30.83\pm 11.43$
w/ image masking (default) $68.50\pm 1.78$ $68.00\pm 6.18$ $77.83\pm 0.82$ $39.33\pm 15.80$
TABLE III: Ablation study on the policy architecture. We explore the effect of the tracks fed into the policy in two positions: transformer input (early fusion) and MLP head (late fusion), as illustrated in Figure 3.
early fusion late fusion Spatial Object Goal Long
$44.67\pm 1.84$ $56.67\pm 3.09$ $5.33\pm 0.24$ $22.33\pm 4.94$
$65.50\pm 3.89$ $60.00\pm 1.47$ $72.83\pm 4.73$ $42.76\pm 14.62$
$68.50\pm 1.78$ $68.00\pm 6.18$ $77.83\pm 0.82$ $39.33\pm 15.80$
Effect of track length. The track length of our pre-trained track transformer is 16 and we utilize all 16 points of each track by default as input for our policy. In this experiment, we investigate the effect of track length of our policy input, by training our policy with the predicted tracks truncated into 4, 8, and 16 steps. The results in Figure 9 demonstrate that longer tracks result in higher success rates in general, except LIBERO-Object where track length $=8$ yields the best performance. In particular, reducing track length leads to larger performance drops on LIBERO-Goal and LIBERO-Long. These tasks require a comprehensive understanding of the task goal. Longer tracks provide more detailed and specific subgoals for each task, which is crucial for guiding the agent’s movements and actions in subsequent stage. Conversely, for LIBERO-Object, which emphasizes precision operations over understanding of the task goal, a policy with a length of 16 underperforms slightly compared to the model with a length of 8. We hypothesize that the longer tracks might interfere with the learning of inverse dynamics due to noise. This also supports our approach of employing tracks both as subgoal conditions and in leveraging the nature of inverse dynamics.
Effect of image masking in track transformer. When training the track transformer, we randomly mask out patches in the images and learn to reconstruct them as an auxiliary task. We conduct an ablation study by removing the image masking loss and comparing it against our standard configuration, where we randomly mask out $50\%$ of the image masks. In Table II, the results reveal a slight decline in policy performance when image masking is omitted, suggesting image masking can be a useful auxiliary task.
Effect of early and late fusion in policy architecture. As shown in Figure 3, the predicted tracks are fed into the policy both before and after the transformer architecture within the policy, which we call early fusion and late fusion respectively. We conduct ablation studies by removing these two track inputs. The results are shown in Table III. We can see that removing the late fusion leads to the most significant performance drop; on LIBERO-Goal, w/ only early fusion performs similarly to other baselines, whereas only late fusion performs marginally worse than our full method. This suggests a late fusion of the predicted tracks acts as useful subgoals that help the policy better understand the tasks in a multi-task learning setting. The subgoal prediction is more robust as it is trained on a larger video dataset.
VI Conclusions
In this work, we present an any-point trajectory modeling framework as video pre-training that effectively learns behaviors and dynamics from action-free video datasets. After pre-training, by learning a track-guided policy, we demonstrate significant improvements over prior state-of-the-art approaches and show effective learning from out-of-distribution human videos. We show that a particle-based representation is interpretable, structured, and naturally incorporates physical inductive biases such as object permanence. We hope our works will open doors to more exciting directions in learning from videos with structured representations.
One limitation is that our approach still relies on a set of action-labeled demonstration trajectories for mapping to actions, which limits the generalization of the learned policies. Future works can consider learning the trajectory-following policies using reinforcement learning so that no additional demonstration data are needed. Another limitation of our method is that the video dataset we use in this paper only contains small domain gaps. Learning from in-the-wild video dataset poses additional challenges such as multi-modal distribution, diverse camera motions, and sub-optimal motions. We leave these extensions for future work.
VII Acknowledgement
This work was supported in part by the BAIR Industrial Consortium, and the InnoHK of the Government of the Hong Kong Special Administrative Region via the Hong Kong Centre for Logistics Robotics. This work is also supported by the Ministry of Science and Technology of the People’s Republic of China, the 2030 Innovation Megaprojects “Program on New Generation Artificial Intelligence” (Grant No. 2021AAA0150000), and the National Key R&D Program of China (2022ZD0161700). This work is done when Chuan Wen is visiting UC Berkeley.
Appendix A Additional Experimental Results
A-A Simulation Experiments
[…]
Attention map visualization. To demonstrate how tracks guide the policy, we visualize the attention maps between the spatial CLS token and RGB tokens in the spatial transformer of BC and our method. Figure 10 demonstrates that our method effectively focuses on the relevant spatial regions, as specified by the textual instructions. Specifically, it attends to the cream cheese, bowl, and wine bottle in the respective example tasks, while BC is usually distracted by the irrelevant regions, highlighting the superior capability of the tracks as better task prompts.
A-B Discussions on the UniPi Baselines
UniPi [11] proposes to train a language-conditioned video diffusion model $f_{\theta}$ during video pre-training. During policy learning, given the initial image observation $o_{0}$ and the language $l$, UniPi first generates all future frames $\tilde{o}_{1},\dots\tilde{o}_{T-1}=f_{\theta}(o_{0},l)$ and then learns an inverse dynamics model that predicts the action at each time step $a_{t}=\pi(o_{t},\tilde{o}_{t+1})$. UniPi then executes the actions open-loop. However, training a diffusion model to predict the full video can be computation intensive, As such, the UniPi implementation in our paper follows the one in Ko et al. [21], where we predict $N=7$ future frames as the sub-goals for the policy, denoted as $\tilde{o}_{1},\dots\tilde{o}_{N}$. During training, we evenly sample $N$ frames in an episode for training the video prediction model. During policy learning, we train a goal-conditioned policy $\pi(a_{t},\text{done}|o_{t},\tilde{o}_{i})$, where $i\in{1,\dots N}$ is the image sub-goal and $t$ denotes the current timestep. The policy additionally predicts a done flag to determine when it should switch from the current sub-goal $\tilde{o}_{i}$ to the next sub-goal $\tilde{o}_{i+1}$. ATM’s superiority over UniPi can be attributed to two reasons. First, ATM uses a more structured sub-goal representation of point trajectories. Second, ATM performs closed-loop inference, proposing a new sub-goal at each time step, while UniPi’s video diffusion process is too slow to be referenced at every time step.
TABLE IV: Computation cost for performing a single task consisting of 600 time steps for different methods. ATM performs trajectory generation at each time step while being computationally efficient. UniPi performs an open-loop future goal generation once at the beginning using a video diffusion model. UniPi-Replan generates image sub-goals at more frequent time steps and is the slowest.
Method ATM UniPi UniPi-Replan
Computation (TFLOPS) 1.56 39.29 2946.75
To train the UniPi policy, we sample $o_{i}$ from a future frame $o_{t}$ where $t\in[t,t+t_{max}]$. We choose $t_{max}$ for each task suite ($t_{max}=16$ for LIBERO-Object and LIBERO-Spatial, $t_{max}=50$ for LIBERO-Goal and LIBERO-Long). To mitigate dataset imbalance when learning the done flag, we sample $o_{i}$ as the next frame 10% of the time. We perform MSE regression on both the action $a_{t}$ and done.
In order to decouple the two advantages of ATM over UniPi additionally compare with a UniPi variation where we train the video prediction model to predict a fixed time step into the future $\tilde{o}_{t+H}=f_{\theta}(o_{t})$ for every $H$ steps of policy execution, where $H=8$. As this variation replans the sub-goal more frequently, we call this method UniPi-Replan, similar to the implementation in [5]. The results are shown in Table V. Surprisingly, we found that this variation performs even worse than UniPi. We thus draw the conclusion that a structured sub-goal representation can be much more effective than an image sub-goal. The reason is that predicting an image goal at a fixed future time step can be a difficult objective for the video prediction model, leading to inconsistent subgoals. Please refer to the failure videos of various baselines on our website. Additionally, this method requires heavy computation. A comparison of the computation needed is shown in Table IV. Due to the computation cost, we only evaluate UniPi-Replan on LIBERO-Object and LIBERO-Goal and report the average success across 10 trials on one policy training random seed.
Figure 10: The attention maps of BC and Ours in the spatial transformer. We extract the attention weights between spatial CLS tokens and RGB tokens, highlighting the policy’s focus on specific spatial regions during decision-making. The heatmaps reveal our policy’s targeted attention on task-relevant areas, in contrast to BC’s tendency to focus on irrelevant backgrounds. This underscores the effectiveness of input tracks in the spatial transformer as good task prompts, guiding the CLS token to attend to appropriate areas.
TABLE V: Average success rate on LIBERO benchmark. Our method performs consistently better than all the baselines across all suites. UniPi-Replan is only evaluated for a single seed due to the computation cost.
Method Libero-Spatial Libero-Object Libero-Goal Libero-Long Libero-90
BC-Full-Trainset (Oracle) $71.83\pm 3.70$ $71.00\pm 7.97$ $76.33\pm 1.31$ $24.17\pm 2.59$
BC $39.00\pm 8.20$ $51.83\pm 13.54$ $42.50\pm 4.95$ $16.67\pm 3.66$ $29.78\pm 1.14$
R3M-finetune [31] $49.17\pm 3.79$ $52.83\pm 2.25$ $5.33\pm 1.43$ $9.17\pm 2.66$ $9.59\pm 0.27$
VPT [3] $37.83\pm 4.29$ $19.50\pm 0.82$ $3.33\pm 2.36$ $3.83\pm 1.65$
UniPi [11, 21] $\bf 69.17\pm 3.75$ $59.83\pm 3.01$ $11.83\pm 2.02$ $5.83\pm 2.08$
UniPi-Replan [5] - 31.00 3.00 - -
ATM (Ours) $68.50\pm 1.78$ $\bf 68.00\pm 6.18$ $\bf 77.83\pm 0.82$ $\bf 39.33\pm 15.80$ $\bf 48.41\pm 2.09$
Figure 11: To summarize spatial information, we perform self-attention on a sequence consisting of all views’ track and image patches and a CLS token. To integrate information across time, we perform casual self-attention between spatial CLS token, proprioception, and an action CLS token per timestep. To regress actions, we concatenate each timestep’s action CLS token and proposed tracks.
Figure 12: Given a video (left), we query 1000 randomly sampled points using an off-the-shelf TAP model (middle), where each colored dot represents the starting position of a track. We then filter the tracks using a heuristic of their position displacement across the video and re-sample around these points (right). We can see that extracted tracks are concentrated around informative objects, such as the robot’s gripper and manipulation targets.
A-C Human-to-robot Transfer Details
To demonstrate the potential of ATM to leverage out-of-domain videos, we design human-to-robot transfer tasks in three different settings: 1) deformable object: fold the cloth and pull it to the right, 2) long horizon: put the tomato into the pan and close the cabinet door, and 3) tool using: use the broom to sweep the toys into the dustpan and put it in front of the dustpan. We collect 10 robot teleportation trajectories (action-labeled) and 100 human manipulation videos (action-free). We compare three methods: 1) behavioral cloning only with 10 action-labeled robot demos, 2) ATM, training track transformer with only 10 robot demos, and 3) ATM, training a track transformer with both action-free human and action-labeled robot data. For each task, we train each method with three different random seeds, evaluate them in the real world 10 times, and report the average success rate in Table I.
Human-to-Robot visualization. The detailed video visualizations of human-to-robot skill transfer are shown in Figure 13. Due to limited training samples, the track transformer trained with only 10 robot videos fails to predict the future point trajectories, leading to low success rates. In contrast, the trajectory model trained with human videos generates high-quality tracks in each frame, guiding the agent to successfully complete the task with only 10 action-labeled trajectories. This indicates our any-point trajectory modeling enables to transfer the motion prior from cross-embodiment videos to robot skills, which significantly improves policy learning.
TABLE VI: Hyperparameters of track transformer training.
TABLE VII: Hyperparameters of policy training.
Hyperparameters Track Transformer
epoch 100
batch size 1024
optimizer AdamW
learning rate 1e-4
weight decay 1e-4
lr scheduler Cosine
 lr warm up 5
clip grad 10
point sampling variance filtering
number of points 32
track length 16
track patch size 4
image mask ratio 0.5
augmentation ColorJitter,RandomShift
Hyperparameters Policy
epoch 100
batch size 512
optimizer AdamW
learning rate 5e-4
weight decay 1e-4
lr scheduler Cosine
lr warm up 0
clip grad 100
point sampling grid
number of points 32
track length 16
frame stack 10
augmentation ColorJitter,RandomShift
TABLE VII: Hyperparameters of policy training.
Figure 13: The visualizations of human demos and rollout videos of ATM policies trained with and without human data. We can see that ATM is able to take advantage of out-of-domain videos, i.e., human videos, to generate more precise tracks, resulting in better policy performance.
Appendix B Implementation Details
B-A Policy Architecture
We include a more detailed visual of the ViT-T [25, 19] policy used in our experiments in Figure 11. The input to our policy is a set of temporally stacked images across multiple views $o_{t}\in\mathbb{R}^{V\times T\times C\times H\times W}$, and proprioception $p_{t}\in\mathbb{R}^{D_{p}}$. To process these inputs, ViT-T consists of three stages:
Spatial Encoding: We encode all modalities at each timestep. We first leverage the frozen track transformer to propose a set of tracks for all $V$ frames in $o_{t}$. We then project each modality with modality-specific encoders into a shared embedding space $\mathbb{R}^{D}$. Each modality’s tokens are embedded using a shared learned modality token, and modality-specific positional embeddings. We then concatenate tokens across all modalities and views with a learned spatial CLS token, and perform self-attention on the sequence. We extract the spatial CLS token as the representation.
Temporal Decoding: We process the encoded modalities across timesteps into actions. We first project the proprioception at each timestep into the same shared embedding space $\mathbb{R}^{D}$. We then interleave the encoded proprioception, the spatial CLS token, and a learned action CLS token across timesteps into a sequence, before performing causally-masked self attention between the sequence.
Action Head: We treat each timestep’s output independently and parameterize actions using an MLP. For each timestep, we take the action CLS token, and fuse the CLS token with the reconstructed tracks of the current timestep.
B-B Efficient Training with Point Filtering
We adopt a heuristic to filter out the static points in the background and then utilize an off-the-shelf tracking model to generate the corresponding tracks of the large-motion points. The visualization of the sampled points before and after the filtering process is shown in Figure 12.
B-C Training Details
We list the training hyperparameters for the track transformer and track-guided policy in Table VII and VII, which are fixed for all experiments on LIBERO benchmark. We train all models on 4 A100 GPUs with DeepSpeed strategy.
We train the track transformer using the ground truth tracks generated by CoTracker [18] and save the checkpoint with the lowest validation loss as our final model to apply for policy learning. We do not incorporate frame stacking for track transformer to avoid causal confusion [48].
We train policies using the expert demonstrations provided by LIBERO, which is collected by human experts through teleoperation with 3Dconnexion Spacemouse [25]. Since the validation loss in behavioral cloning is not always reliable, we save the checkpoint of the last epoch for online rollout evaluation.